Lovell Computer Parts Inc. is in the process of setting a selling price on a new component it has just designed and developed. The following cost estimates for this new component have been provided by the accounting department for a budgeted volume of 50,000 units.

Per Unit Total
Direct materials $52
Direct labor $22
Variable manufacturing overhead $18
Fixed manufacturing overhead $650,000
Variable selling and administrative expenses $16
Fixed selling and administrative expenses $400,000


Lovell Computer Parts management requests that the total cost per unit be used in cost-plus pricing its products. On this particular product, management also directs that the target price be set to provide a 30% return on investment (ROI) on invested assets of $1,000,000.Compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 30% on this new component. (Round markup percentage to 2 decimal places, e.g. 10.50%.)

Markup percentage %
Target selling price $

eTextbook and Media

Assuming that the volume is 40,000 units, compute the markup percentage and target selling price that will allow Lovell Computer Parts to earn its desired ROI of 30% on this new component. (Round answers to 2 decimal places, e.g. 10.50% or 10.50.)

Markup percentage %
Target selling price $

eTextbook and Media
Answer #1

a) Variable cost per unit= Direct materials+Direct labor+Variable manufacturing overhead+Variable selling and administrative expenses

= $52+22+18+16= $108

Fixed cost per unit= (Fixed manufacturing overhead+Fixed selling and administrative expenses)/Units

= $(650000+400000)/50000

= $21

Total cost per unit= Variable cost per unit+Fixed cost per unit

= $108+21= $129

Expected income= Invested assets*ROI

= $1000000*30%= $300000

Expected income per unit= $300000/50000= $6

Markup percentage= Expected profit per unit*100/Total cost

= $6*100/129= 4.65%

Target selling price= Total cost per unit+Expected profit per unit

= $129+6= $135

Markup percentage 4.65 %
Target selling price $135

b) Variable cost per unit= Direct materials+Direct labor+Variable manufacturing overhead+Variable selling and administrative expenses

= $52+22+18+16= $108

Fixed cost per unit= (Fixed manufacturing overhead+Fixed selling and administrative expenses)/Units

= $(650000+400000)/40000

= $26.25

Total cost per unit= Variable cost per unit+Fixed cost per unit

= $108+26.25= $134.25

Expected income= Invested assets*ROI

= $1000000*30%= $300000

Expected income per unit= $300000/40000= $7.5

Markup percentage= Expected profit per unit*100/Total cost

= $7.5*100/134.25= 5.59%

Target selling price= Total cost per unit+Expected profit per unit

= $134.25+7.5= $141.75

Markup percentage 5.59 %
Target selling price $141.75